# Quantum Geometry: Revisiting electronic scales in quantum matter

### v1

@@ III Quantum Geometry in Band Theory
In contrast to the Berry curvature, the quantum metric has remained largely unexplored, primarily due to the lack of an intuitive understanding. Moreover, challenges inherent in its definition create additional conceptual hurdles that complicate its interpretation and practical application. The band resolved QGT Eq. (5) corresponds to dipole transitions between the band $n$ and all other bands $m\neq n$, filled or empty, at a given point in the BZ. It is thus qualitatively different from the metric introduced by Kohn, where the projector in Eq. (3) is the ground state projector $\hat{P}_{0}$, see Fig.2. For band electrons, we can define $\hat{P}_{0}$ into all occupied single-particle states as
$\hat{P}_{0}=\int_{\rm BZ}\sum_{n}\Theta(\mu-\varepsilon_{n{\bm{k}}})|u_{n{\bm{k}}}\rangle\langle u_{n{\bm{k}}}|$ (6)
@@
with $\mu$ the chemical potential. In particular, in the case of multiple filled bands, the ground state geometric tensor cannot be obtained by summing $\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})$ over filled bands and momenta. This is because, unlike Berry curvature, the quantum metric is not band additive [31, 32].
@@
The additive property can be explained via Wannier functions [33]. The integral of the quantum metric over the Brillouin zone (BZ) for a single band, $\int_{\rm BZ}{\rm Re}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]=g^{n}_{\mu\nu}$, sets a lower bound on the spatial spread of the Wannier function [34]. When additional bands are included in the construction of the Wannier function, it is reasonable to expect a more localized Wannier state with a reduced spatial spread. However, since the quantum metric is inherently a positive quantity, the sum of the metric of individual bands within a projected manifold cannot be less than the metric of the full manifold. Thus metric cannot be additive, $g_{\mu\nu}\neq\sum_{n\in{\rm occ.}}g^{n}_{\mu\nu}$.
@@
Setting these subtleties aside, the quantum metric and Berry curvature have a close relationship that becomes evident in the case of Landau levels. As we know, the Chern number arises from the non-commutativity of the projected position operators, $\hat{X}=\hat{P}\hat{x}\hat{P}$ and $\hat{Y}=\hat{P}\hat{y}\hat{P}$, with $[\hat{X},\hat{Y}]=i\ell_{B}^{2}$, where $\ell_{B}$ is the magnetic length. The non-commutativity of these operators naturally implies that the variance $\langle\hat{X}^{2}+\hat{Y}^{2}\rangle$ is subject to a lower bound. Remarkably, the lowest Landau level satisfies the so-called trace condition [35], in which this variance saturates the bound. In simpler terms, the cyclotron orbits of an electron in a Landau level have the smallest spatial spread capable of supporting the necessary winding.
Figure 3: Separation of scales in the tight-binding approximation and the emergence of a geometric scale. Panel a. shows the band structure for a free particle in a honeycomb lattice potential. Panel b. shows the optical conductivity.
@@
Although real materials deviate significantly from the idealized limit of Landau levels, the connection between topology and geometry remains true. Topological bands inherently require extended Wannier states, as the quantum metric is subject to a lower bound imposed by the topological index. Thus topology is a sufficient condition for non-trivial quantum geometry. This statement can be made more precise through an estimate of the quantum metric derived from the dielectric constant. Within the space of all insulators, it has been shown that the quantum metric in topological insulators is significantly larger than that in trivial insulators with comparable band gaps [36].
Importantly, topology is sufficient but not necessary for quantum geometry in real materials. Unlike the Berry curvature, the metric is not limited to systems that break time-reversal or inversion symmetries. Quantum geometry is ubiquitous to all known materials. The key challenge lies in developing robust methods to extract it and quantify its impact on the observed behavior.
IV Where to find quantum metric?
@@
The experimental challenge in measuring the metric is two-fold. First, the probe must couple to the symmetric part of the QGT, and second, it must identify the specific geometric scale being probed. This issue ties back to the freedom in the choice of the projector $\hat{P}$ and the complementary projector $1-\hat{P}$ in the definition of the QGT in Eq. (3). Thus, the geometric tensor should not be viewed as an intrinsic property of a band or the ground state but rather as a property of the choice of projected states. These states are determined by the experimental probe, which itself imposes an effective truncation of the Hilbert space. That is, it is natural for the experimental probe to select the appropriate scale. A probe operating at some frequency range $\Delta\omega_{\rm ext}$ will carry an intrinsic time uncertainty $\Delta t\sim 1/\Delta\omega_{\rm ext}$. Alternatively, the Coulomb interaction strength $U$ can itself impose the energy window and which states can be truncated out of the Hilbert space.
@@
IV.1 Instantaneous response and sum rules
@@
The magnitude of dipole fluctuations in a real material depends on the time scale. In a metal, the dynamics of a single electron is diffusive in the long-time limit $t>\tau$ with $\langle r(t)r(0)\rangle\sim Dt$ with $D$ the Drude weight dictated by the band effective mass and carrier density. This diffusive propagation of fluctuations saturates at the mean free path, $\ell_{\rm mr}^{2}\sim D\tau$. Note for momentum preserving, clean metals $\ell_{\rm mr}\to\infty$. On the other hand, the very short-time dynamics $t<\hbar/V$ is governed by the local quantum chemistry of the atom. Quantum geometry introduces a length scale that lies in between these well-understood scales. It is comparable to the bandwidth of the optically active bands and is relevant for dynamics at times shorter than $\tau$ but larger than $\hbar/V$. We can formalize this intuition by rewriting the Souza-Wilkens-Martin (SWM) sum rule in terms of a Fermi surface (Drude) contribution $g_{\rm FS}$, a quantum geometric part due to sharing electrons across multiple orbitals $\ell_{g}^{2}$, and an atomic contribution $\ell_{a}^{2}$,
[…]
@@
where $\sigma_{\mu\mu}(\omega)$ is the longitudinal optical conductivity along one spatial direction. We have introduced a normalization with respect to charge density $n$ to make a direct comparison of this scale of dipole fluctuations to the average inter-atomic distance $a\equiv n^{-1/d}$. This gives units of length squared to the integral at every dimension. The Fermi surface contribution, $g_{\rm FS}$, is either infinite or zero, depending on whether a Fermi surface is present or absent [37]. There is no such divergence for insulators, and the largest scale of dipole fluctuations is controlled by $\ell_{g}$. Fluctuations also exist at a smaller scale, $\ell_{a}$, corresponding to atomic orbitals. Capturing these requires integrating the optical conductivity up to a large energy cutoff $\Omega$ approximately given by the atomic level gap, to remove the contributions of core electron excitations. These high-energy transitions and the corresponding $\ell_{a}\sim a_{B}$ fluctuations are neglected in tight-binding approximations.
@@
At intermediate frequencies between $\omega\in[\tau^{-1},\Omega]$, we can identify the geometric contribution to the localization tensor of the ground state $\ell_{g}$. This contribution is dominated by the band geometry without taking into account filling. The spectral weight at these frequencies would be pushed smoothly into the Drude peak and disappear if we adiabatically take electrons apart into the atomic limit, which means that it appears from the quantum interference and delocalization of orbitals across lattice sites. It can be viewed as the size of “molecular orbitals”. Importantly,the inter-band optical absorption yields a finite value of $\ell_{g}$, both in insulators and in metals. The emergence of this scale can be seen in Fig. 3, where we show the optical absorption for a free particle subjected to a honeycomb lattice potential. The finite energy contribution to optical conductivity, centered at $\omega_{g}$, arises from the quantum interference between the two sublattices, implying that $\ell_{g}$ is comparable to the atomic distances, and the energy $\omega_{g}$ comparable to the hopping energy between sublattices. Applying Eq. (7) to free electrons under a magnetic field of magnitude $B$, we find that $\omega_{g}=\omega_{c}=\hbar/m\ell_{B}^{2}$ and $\ell^{2}_{g}=\ell^{2}_{B}=\hbar/eB$.
@@
Sum rules reveal that polarization fluctuations are responsible for other defining characteristics of bound electrons, such as their optical mass $\hbar/m\sim\langle{\bm{r}}\cdot{\bm{v}}\rangle$, part of orbital magnetization $M\sim\langle{\bm{r}}\times{\bm{v}}\rangle$, or shot noise $\langle{\bm{v}}\cdot{\bm{v}}\rangle$. As outlined in [38], the generalized sum rules can be systematically obtained from different time derivatives of the unequal time dipole-dipole correlator, or time-dependent QGT, $\mathcal{Q}_{\mu\nu}(t)=\langle\hat{r}_{\mu}(t)(1-\hat{P})\hat{r}_{\nu}\rangle$. Taking these derivatives amounts to convolute the geometric tensor with various powers of the energy differences at which the dipole transitions occur, generalizing the SWM sum rule one finds,
[…]
@@
In the right-hand side band indices $n,m$ are summed over filled and empty states respectively, and the momentum dependence is implicit. In the left hand side the dissipative conductivity contains both the longitudinal and Hall components which pick up the symmetric and and antisymmetric parts of the geometric tensor. We introduce a low energy cutoff to remove any Fermi surface contribution. Assuming that the geometric contribution to the optical conductivity is dominated by a characteristic energy $\omega_{g}$, the optical mass may be approximated by $1/m\sim\omega_{g}\ell_{g}^{2}$, the shot noise by $\omega_{g}^{2}\ell_{g}^{2}$ and the dielectric constant of an insulator $\varepsilon=1+\chi$ with $\chi\sim\ell_{g}^{2}/\omega_{g}$. Note that this also highlights that moments with positive powers of energy differences $\omega^{\eta}$, $\eta>0$ become hard to estimate without accounting for the atomic contribution at high energies $\sim\omega_{a}^{\eta}\ell_{a}^{2}$. While this reasoning applies perfectly in a system with flat bands, such as in the Landau level problem, strictly speaking all different sum rules, or moments of ground state dipole fluctuations, should be computed by the knowledge of all states in the Hilbert space, both filled and empty, and this approximation may not be suited for highly dispersive bands and broad distributions of $\sigma(\omega)$. In this limit, we can instead keep track of these few moments as independent geometric scales, which can be used to accurately describe the response of materials beyond the semiclassical regime.
@@
Other than estimations of geometric quantities, optical sum rules have been used to establish rigorous bounds on localization [39, 40, 41, 42], superfluid stiffness [43, 44, 45, 46], dielectric function [36, 47], and more recently, on the energy gap in topological insulators[48].
@@
Geometric quantities also appear in density response functions at small $q$, as the density and position operators satisfy $\hat{\rho}_{q}\approx 1+iq_{\mu}\hat{r}_{\mu}$. This relation was recently employed to generalize sum rules to non-linear order using the density operator [49]. Furthermore, it provides a pathway to extract the QGT from the static structure factor [50, 51], which can then be utilized to establish topological bounds [52]. While such sum rules and related bounds are excellent for building intuition, they are challenging to measure or verify experimentally. Although some progress has been made [53, 54, 55], there is increasing evidence that extracting the quantum metric via sum rules in principle requires a sum over all frequencies [32].
IV.2 Transport coefficients
@@
The quantum geometry of the ground state is an essential ingredient for the characterization of a bulk quantum phase. Nevertheless, its observational status remains rather complicated. These difficulties are intimately linked to its definition via the overlap between the projector $\hat{P}$ and the complement $1-\hat{P}$ (Eq. (3)), which means one should probe the system adiabatically, while simultaneously keeping track of all other energies. This somewhat paradoxical protocol is by all accounts hard to achieve but not impossible [56]. To give some orientation, in the following we will elucidate the main factors which enter into the integrand of a response function. However, to avoid an overly technical discussion, we will keep a loose definition of $\ell_{g}$, remembering that depending on the response, different and linearly independent geometric moments or even combinations of moments may contribute [57] which are usually tensorial objects and may differ from the definition in Eq. (7). We present in Table 1 the exact forms and references regarding these contributions.
@@
Berry curvature matrix elements $\Omega_{\mu\nu}^{nm}$ $\ell_{g}^{2}$ $(r_{\mu}^{nm}r_{\nu}^{mn}-r_{\nu}^{nm}r_{\mu}^{mn})$
Quantum metric matrix elements $g^{nm}_{\mu\nu}$ $\ell_{g}^{2}$ $(r_{\mu}^{nm}r_{\nu}^{mn}+r_{\nu}^{nm}r_{\mu}^{mn})$
band resolved Berry curvature $\Omega_{\mu\nu}^{n}$ $\ell_{g}^{2}$ $\sum_{m\neq n}\Omega_{\mu\nu}^{nm}$
band resolved quantum metric $g^{n}_{\mu\nu}$ $\ell_{g}^{2}$ $\sum_{m\neq n}g_{\mu\nu}^{nm}$
Landau Zener coupling $G_{\mu\nu}^{n}$ $\ell_{g}^{2}/\omega_{g}$ $\sum_{m\neq n}g_{\mu\nu}^{nm}/\omega_{nm}$
Quantum connection $Q^{nm}_{\mu\nu\lambda}$ $\ell_{g}^{3}$ $r^{nm}_{\mu}r^{mn}_{\nu}(r^{mm}_{\lambda}-r^{nn}_{\lambda}+i\partial_{\lambda}\operatorname{log}r_{mn}^{\mu})$
@@
Anomalous Hall $\sigma_{\mu\nu}$ $\ell_{g}^{2}$ $\frac{e^{2}}{\hbar}\int_{\rm BZ}\sum_{n}f_{n}\Omega^{n}_{\mu\nu}$ [58, 59, 8, 60]
@@
Nonlinear anom. Hall $\sigma^{(2)}_{\mu\mu\nu}$ $\ell_{\rm mr}\ell_{g}^{3}$ $\tau\frac{e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum_{n}f_{n}\partial_{\mu}\Omega^{n}_{\mu\nu}$ [61, 62, 63]
Static susceptibility $\chi_{\mu\nu}$ $\ell_{g}^{2}/\omega_{g}$ $\frac{e^{2}}{2\hbar}\int_{\rm BZ}\sum_{n}f_{n}G^{n}_{\mu\nu}$ [36]

### v2

@@ Introduction
In contrast, the quantum metric has remained largely unexplored, primarily due to the lack of an intuitive understanding. Its effects on semiclassical equations of motions are only now being explored [36]. Moreover, challenges inherent in its definition create additional conceptual hurdles that complicate its interpretation and practical application. The band resolved QGT (Eq. (5)) corresponds to dipole transitions between the band $n$ and all other bands $m\neq n$, filled or empty, at a given point in the BZ. It is thus qualitatively different from the metric introduced by Kohn, where the projector in Eq. (3) is the ground state projector $\hat{P}^{0}$ (Fig.2). For band electrons, we can define $\hat{P}^{0}$ into all occupied single-particle states as
$\hat{P}^{0}=\int_{\rm BZ}\sum_{n}\Theta(\mu-\varepsilon_{n{\bm{k}}})|u_{n{\bm{k}}}\rangle\langle u_{n{\bm{k}}}|$ (6)
@@
with $\mu$ the chemical potential. In particular, in the case of multiple filled bands, the ground state geometric tensor cannot be obtained by summing $\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})$ over filled bands and momenta. This is because, unlike Berry curvature, the quantum metric is not band additive [37].
Figure 3: Separation of scales and the emergence of a geometric length scale from lattice interference. a. Band structure for a free particle in a periodic honeycomb lattice potential. The tight-binding approximation is valid when the potential $V$ is deep enough to spectrally isolate a few orbitals, here the two lower connected bands. The arrows indicate optical transitions within a single band (orange), within the low-energy band manifold (blue), and between the low- and high-energy bands (black). b. Optical conductivity for the same system, where low-frequency spectral weight features a Drude weight due to the Fermi surface (orange) in addition to a geometric spectral weight (blue), due to optical transitions between the same orbital across distinct lattice sites. The higher-frequency features reflect dipole transitions that correspond to fluctuations at the scale of an atomic orbital, truncated away in the tight-binding approximation. Looking at the spectral weight at low energies (blue) we can define an effective geometric length scale, $\ell_{g}$, that governs dipole fluctuations due to lattice interference, as well as a resonant energy $\mathcal{E}$ associated with its local dynamics. The resonant energy $\mathcal{E}$ is defined by a weighted average of band energy differences $\varepsilon_{n{\bm{k}}}-\varepsilon_{m{\bm{k}}}$ by geometric factors and can often be approximated by the first dominant peak in optical conductivity, comparable to the hopping energy. The boxes indicate the characteristic length scales associated with the dipole fluctuations contributing to the long-wavelength (orange), intermediate (blue), and atomic-scale (black) spectral weight.
@@
The fact that the quantum metric is not additive can be explained via Wannier functions. The integral of the quantum metric over the Brillouin zone (BZ) for a single band, $\int_{\rm BZ}{\rm Re}[\mathcal{Q}^{n}_{\mu\nu}({\bm{k}})]=g^{n}_{\mu\nu}$, sets a lower bound on the spatial spread of the Wannier function [38, 39]. When additional bands are included in the construction of the Wannier function, it is reasonable to expect a more localized Wannier state with a reduced spatial spread. However, since the quantum metric is inherently a positive quantity, the sum of the metric of individual bands within a projected manifold cannot be less than the metric of the full manifold. Thus the metric cannot be additive, $g_{\mu\nu}\neq\sum_{n\in{\rm occ.}}g^{n}_{\mu\nu}$. The metric is consequently not a conventional response function and depends on both filled and unfilled bands [40]. It is important to distinguish this measure of spread from exponential localization; for instance, in topological bands, Wannier functions are obstructed and exhibit power-law tails, even though their spatial spread as measured by the quantum metric remains finite. Therefore, also for topological bands the geometric scales introduced by dipole fluctuations are finite and well defined.
@@
Setting these subtleties aside, the quantum metric and Berry curvature have a close relationship that becomes evident in the case of Landau levels. As we know, the Chern number arises from the non-commutativity of the projected position operators in the lowest Landau level, $\hat{X}=\hat{P}\hat{x}\hat{P}$ and $\hat{Y}=\hat{P}\hat{y}\hat{P}$, with $\langle[\hat{X},\hat{Y}]\rangle=i\ell_{B}^{2}$, where $\ell_{B}$ is the magnetic length. The non-commutativity of these operators naturally implies that the variance $\langle\hat{X}^{2}+\hat{Y}^{2}\rangle$ is subject to a lower bound. Remarkably, the lowest Landau level satisfies the so-called trace condition [28], in which this variance saturates the bound. In simpler terms, the cyclotron orbits of an electron in a Landau level have the smallest spatial spread capable of supporting the necessary winding.
@@
Although real materials deviate significantly from the idealized limit of Landau levels, the connection between topology and geometry remains true. Topological bands inherently require extended Wannier states, as the quantum metric is subject to a lower bound imposed by the topological index. Thus topology is a sufficient condition for non-trivial quantum geometry. This statement can be made more precise through an estimate of the quantum metric derived from the dielectric constant. Within the space of all insulators, it has been shown that the quantum metric in topological insulators is significantly larger than that in trivial insulators with comparable band gaps [41]. Importantly, topology is sufficient but not necessary for quantum geometry in real materials. Unlike the Berry curvature, the quantum metric is not restricted to systems that break time-reversal or inversion symmetry. Quantum geometry is present in all known materials in various amounts, and the key challenge is to quantify its influence on physical observables (Box 1).
Where to find the quantum metric?
@@
The experimental challenge in measuring the metric is two-fold. First, the probe must couple to the symmetric part of the QGT, and second, it must identify the specific geometric scale being probed. This issue ties back to the freedom in the choice of the projector $\hat{P}$ and the complementary projector $1-\hat{P}$ in the definition of the QGT in Eq. (3). Thus, the geometric tensor should not be viewed as an intrinsic property of a band or the ground state but rather as a property of the choice of projected states. These states are determined by the experimental probe, which itself imposes an effective truncation of the Hilbert space. That is, it is natural for the experimental probe to select the appropriate scale. A probe operating at a frequency range $\Delta\omega$ carries an intrinsic time uncertainty $\Delta t\approx 1/\Delta\omega$. Alternatively, the Coulomb interaction strength $U$ can impose the energy window, defining the projectors $\hat{P}$ and $1-\hat{P}$ and, therefore, which states are truncated. Unlike the Chern number, which is topologically robust and does not change when trivial bands are truncated, the quantum metric is sensitive to such truncations. Therefore, these choices, and the resulting interpretation of geometric quantities, must be critically considered when approaching a condensed matter problem.
@@
Instantaneous response and sum rules
@@
The magnitude of dipole fluctuations in a real material depends on the time scale. In the long time limit, $t>\tau$, the dynamics of a single electron in a metal is diffusive with $\langle r(t)r(0)\rangle\approx Dt$ where $D$ is the Drude weight dictated by the band effective mass and carrier density. Practically, this diffusive propagation of fluctuations saturates at the mean free path, $\lambda^{2}\approx D\tau$. Note that for momentum preserving, clean metals $\lambda\to\infty$. Conversely, the very short-time dynamics $t<\hbar/V$ is governed by the local quantum chemistry of the atom.
Quantum geometry introduces a length scale that lies in between these well-understood scales. It is comparable to the bandwidth of the optically active bands and is relevant for dynamics at times shorter than $\tau$ but larger than $\hbar/V$. We can formalize this intuition by rewriting the Souza-Wilkens-Martin (SWM) sum rule in terms of a Fermi surface (Drude) contribution $g_{\rm FS}$, and a quantum geometric part due to sharing electrons across multiple orbitals $\ell_{g}^{2}$
$\dfrac{\hbar}{\pi ne^{2}}\int\limits_{0}^{\infty}d\omega\dfrac{{\rm Re}[\sigma_{\mu\mu}(\omega)]}{\omega}={g_{\rm FS}\over n}+\ell_{g}^{2},$ (7)
@@
where $\sigma_{\mu\mu}(\omega)$ is the longitudinal optical conductivity along one spatial direction. We have introduced a normalization with respect to charge density $n$ to make a direct comparison of this scale of dipole fluctuations to the average inter-atomic distance $a\equiv n^{-1/d}$. This gives units of length squared to the integral at every dimension. The Fermi surface contribution, $g_{\rm FS}$, is either infinite or zero, depending on whether a Fermi surface is present or absent [1]. There is no such divergence for insulators, and the largest scale of dipole fluctuations is controlled by the wavefunctions, $\ell_{g}$. Note that even in the absence of lattice interference, fluctuations exist at the smaller scale of the Bohr radius, $\ell_{a}\approx a_{B}$, which define the atomic orbitals. Capturing these requires integrating the optical conductivity up to a large energy cutoff, $V$, approximately given by the atomic level gap, to remove the contributions of core electron excitations. These high-energy transitions and fluctuations are effectively neglected in tight-binding approximations.
@@
At intermediate frequencies between $\omega\in[\tau^{-1},V]$, we can identify the geometric contribution to the localization tensor of the ground state $\ell_{g}$. This contribution is dominated by the band geometry without taking into account filling. The spectral weight at these frequencies is continuously transferred into the Drude peak and vanishes as the system is adiabatically taken to the atomic limit. This implies that it originates from quantum interference and the delocalization of orbitals across lattice sites. It can be viewed as the size of “molecular orbitals”, resonant across multiple sites in the unit cell. Importantly, the interband optical absorption yields a finite value of $\ell_{g}$, both in insulators and in metals. The emergence of this scale is shown in Fig. 3 with optical absorption for a free particle subjected to a honeycomb lattice potential. The finite energy contribution to optical conductivity, centered at $\mathcal{E}$, arises from the quantum interference between the two sublattices, implying that $\ell_{g}$ is comparable to the atomic distances, and the energy $\mathcal{E}$ can be interpreted as a resonance energy comparable to the hopping energy between sublattices. Applying Eq. (7) to free electrons under a magnetic field of magnitude $B$, we find that $\mathcal{E}/\hbar=\hbar/m\ell_{B}^{2}$ is the cyclotron frequency and $\ell^{2}_{g}=\ell^{2}_{B}=\hbar/eB$ the magnetic length. However, this interpretation can be in fact generally applied to multi-orbital materials simply from the fact that bands are generally composed of admixtures of orbitals separated spatially.
@@
Sum rules reveal that polarization fluctuations are responsible for other defining characteristics of bound electrons, such as their optical mass $1/m\approx\langle{\bm{r}}\cdot{\bm{v}}\rangle$, part of orbital magnetization $M\approx\langle{\bm{r}}\times{\bm{v}}\rangle$, or shot noise $\langle{\bm{v}}\cdot{\bm{v}}\rangle$. The generalized sum rules can be systematically obtained [42] from different time derivatives of the unequal time dipole-dipole correlator, or time-dependent QGT, $\mathcal{Q}_{\mu\nu}(t)=\langle\hat{r}_{\mu}(t)(1-\hat{P})\hat{r}_{\nu}\rangle$. The complementary projector $1-\hat{P}$ in $\mathcal{Q}(t)$ guarantees its gauge independence, which highly simplifies the form of response functions such as polarizability or conductivity. Taking time derivatives of $\mathcal{Q}(t)$ amounts to convoluting the geometric tensor with various powers of the energy differences at which the dipole transitions occur, and generalizing the SWM sum rule one finds,
$\int_{0^{+}}^{\infty}d\omega~{\sigma^{\rm dis}_{\mu\nu}(\omega)\over\omega^{\eta+1}}=\dfrac{\pi e^{2}}{\hbar}\int_{\rm BZ}\omega^{\eta}_{mn}\left(g^{mn}_{\mu\nu}-\dfrac{i}{2}\Omega^{mn}_{\mu\nu}\right).$ (8)
On the right-hand side, band indices, $n,m$, are summed over filled and empty states respectively, and the momentum dependence is implicit. On the left hand side, the dissipative conductivity contains both the longitudinal and Hall components which pick up the symmetric and and antisymmetric parts of the geometric tensor. We introduce a low energy cutoff to remove any Fermi surface contribution, separating the finite geometric contribution of various sum rules to a possibly divergent Fermi surface part. From this we can see that if we assume that the geometric contribution to the optical conductivity is dominated by a characteristic frequency $\mathcal{E}/\hbar$, obtained from an weighted average of $\omega_{mn}$, we can obtain a good estimate for various instantaneous properties of insulators. This approximation is closely related to the Penn model, where the entire set of interband transitions is collapsed into a single effective energy scale (the “Penn gap”), chosen to satisfy optical sum rules [43]. Their optical mass may be approximated by $1/m\approx\mathcal{E}\ell_{g}^{2}$, shot noise by $\mathcal{E}^{2}\ell_{g}^{2}$ and the dielectric constant by $\varepsilon=1+\chi$ with $\chi\approx\ell_{g}^{2}/\mathcal{E}$. Larger $\ell_{g}$ implies larger fluctuations at all moments, and it is indeed the reason behind the large refractive index of silicon, diamond or topological insulators [41].
@@
We should also highlight that moments with positive powers of energy differences $\hbar\omega_{mn}^{\eta}$, $\eta>0$ become hard to estimate without accounting for the atomic contribution at high energies. For this reason it is not possible to accurately capture orbital magnetization without involving core orbitals [44], as one example. This approximation becomes exact in a system with only flat bands, such as in the Landau level problem. However, strictly speaking, all moments of the ground state dipole fluctuations should be computed by the knowledge of all states in the Hilbert space, both filled and empty. Particularly for highly dispersive bands and broad distributions of $\sigma(\omega)$, approximating a single length scale $\ell_{g}$ may not be suitable. In such cases, however, we may instead keep track of these few moments as independent geometric scales, which can be used to accurately describe the response of materials beyond the single band and long-wavelength approximations, still compressing the wavefunction data to a few phenomenological numbers.
@@
Other than estimations of geometric quantities, optical sum rules have been used to establish rigorous bounds on localization [45, 46, 47, 48], superfluid stiffness [49, 50, 51, 52, 53], dielectric function [41, 54], optical mass [55] and more recently, on the energy gap in topological insulators [56].
@@
Geometric quantities also appear in density response functions at small $q$, as the density and position operators satisfy $\hat{\rho}_{q}\approx 1+iq_{\mu}\hat{r}_{\mu}$. This relation was recently employed to generalize sum rules to non-linear order using the density operator [57]. Furthermore, it provides a pathway to extract the QGT from the static structure factor [58, 59], which can then be utilized to establish topological bounds [60]. While such sum rules and related bounds are excellent for building intuition, they are challenging to measure or verify experimentally, although some progress has been made [61, 62, 63, 64].
Transport coefficients
The quantum geometry of the ground state is an essential ingredient for the response characterization of a macroscopic quantum state. In the example of diamond and rock salt, which are both insulators with comparable, large band gaps, their optical properties such as their refractive index ($n^{2}\approx\ell_{g}^{2}/\mathcal{E}$) differ significantly due to the scale $\ell_{g}$ being larger in covalent than ionic bonding[41]. Nevertheless, measuring quantum geometry directly remains rather complicated. These difficulties are intimately linked to its definition via the overlap between the projector $\hat{P}$ and the complement $1-\hat{P}$ (Eq. (3)), which means one should probe the system adiabatically, while simultaneously keeping track of all other energies. This somewhat paradoxical protocol is by all accounts hard to achieve but not impossible [65].
@@
To give some orientation, in the following we will elucidate the main factors which enter into the integrand of a response function. However, to avoid an overly technical discussion, we will keep a loose definition of an intrinsic geometric length scale, $\ell$, (we will keep the symbol $\ell_{g}$ to indicate the length scale coming from the symmetric part of the QGT, the quantum metric), remembering that depending on the response, different and linearly independent geometric moments or even, combinations of moments, contribute to such $\ell$ [66] which are usually tensorial objects and may differ from the definition in Eq. (7). We present in Table 1 the exact forms and references regarding these contributions.
Much progress has been made in diagnosing the most well-known part of the QGT, the Berry curvature ($\ell^{2}$), which can be recast as a single projection operation. As mentioned before, the Berry curvature leads to the intrinsic anomalous Hall effect, which describes the transverse, antisymmetric current component that arises in magnetic materials, which thus carry a nonzero momentum average of the Berry curvature. In magnetic insulators, the same mechanism gives rise to the quantum anomalous Hall effect [67, 68]. More recently, high harmonic generation and polarization-sensitive ARPES techniques were able to directly measure the momentum-resolved Berry curvature in SiO2 [69] and WSe2 [70].
Another route to measuring the Berry curvature is to identify observables which couple to its momentum derivatives. For example, the nonlinear intrinsic anomalous Hall effect has been successfully expressed in terms of the quasiparticle lifetime times the momentum derivative of the Berry curvature ($\lambda\ell^{3}$[71, 72, 73]. This derivative has become known as the Berry curvature dipole, which can retain a nonzero momentum average even in the presence of time-reversal symmetry, as long as spatial inversion symmetry is broken. The Berry curvature dipole has become a valuable probe for the Berry curvature in a number of 2d and 3d materials [73] and also offers a particularly enlightening example what ingredients can constitute the geometric length scale $\ell$. Clearly, the derivative of the Berry curvature estimates how violently the Berry curvature changes inside the Brillouin zone, setting a typical momentum space distance $\kappa$ which separates minima and maxima of the Berry curvature. In real space, this translates to a characteristic distance $\kappa^{-1}\rightarrow\ell$ below which a wavepacket remains irrotational. Thus, the Berry curvature dipole constitutes an anomalous component in the equation of motion that is not due to any intrinsic magnetization, but rather due to the tendency of a wavepacket to acquire a finite Hall angle from uncompensated orbital moments in the tail of the envelope function.
@@
Berry curvature matrix elements $\Omega_{\mu\nu}^{nm}$ $\ell^{2}$ $(r_{\mu}^{nm}r_{\nu}^{mn}-r_{\nu}^{nm}r_{\mu}^{mn})$
Quantum metric matrix elements $g^{nm}_{\mu\nu}$ $\ell^{2}$ $(r_{\mu}^{nm}r_{\nu}^{mn}+r_{\nu}^{nm}r_{\mu}^{mn})$
band resolved Berry curvature $\Omega_{\mu\nu}^{n}$ $\ell^{2}$ $\sum_{m\neq n}\Omega_{\mu\nu}^{nm}$
band resolved quantum metric $g^{n}_{\mu\nu}$ $\ell^{2}$ $\sum_{m\neq n}g_{\mu\nu}^{nm}$
Berry curvature dipole $\partial_{\mu}\Omega_{\nu\lambda}^{n}$ $\ell^{3}$ $\sum_{m\neq n}\partial_{\mu}\Omega_{\nu\lambda}^{nm}$
quantum metric dipole $\partial_{\mu}g_{\nu\lambda}^{n}$ $\ell^{3}$ $\sum_{m\neq n}\partial_{\mu}g_{\nu\lambda}^{nm}$
Landau Zener coupling $G_{\mu\nu}^{n}$ $\ell^{2}/\mathcal{E}$ $\sum_{m\neq n}g_{\mu\nu}^{nm}/\omega_{nm}$
Quantum connection $Q^{nm}_{\mu\nu\lambda}$ $\ell^{3}$ $r^{nm}_{\mu}r^{mn}_{\nu}(r^{mm}_{\lambda}-r^{nn}_{\lambda}+i\partial_{\lambda}\operatorname{log}r^{mn}_{\mu})$
Table 1: Examples of geometric quantities. The “Ingredients” column keeps track of various contributions to the response: intrinsic geometric length scales $\ell$ due to position matrix elements and the intrinsic resonance energy $\mathcal{E}$ across bands. $m$ and $n$ label bands and $\hbar\omega_{mn}$ labels the energy difference between bands.
However, no direct measurements of the momentum-resolved quantum metric ($\ell^{2}$) or its momentum average have been reported so far. There have been suggestions that the metric might couple directly to the dc-conductivity [74, 75, 76] in certain fine-tuned settings, but no linear response observable is known which is generically determined solely by the quantum metric. It has also been suggested that the quantum metric dipole ($\ell^{3}$) appears in the nonreciprocal directional dichroism for materials breaking time-reversal symmetry, a nonlocal conductivity component which originates from a coupling to the electric field gradient [77, 78].
@@
Anomalous Hall $\sigma_{\mu\nu}$ $\ell^{2}$ $\frac{e^{2}}{\hbar}\int_{\rm BZ}\sum_{n}f_{n}\Omega^{n}_{\mu\nu}$ [67, 68, 13, 79]
@@
Nonlinear anom. Hall $\sigma^{(2)}_{\mu\mu\nu}$ $\lambda\ell^{3}$ $\tau\frac{e^{3}}{\hbar^{2}}\int_{\rm BZ}\sum_{n}f_{n}\partial_{\mu}\Omega^{n}_{\mu\nu}$ [71, 72, 73]
Non-reciprocal dichroism $\gamma_{\mu\nu\lambda}$ $\ell^{3}$ $e^{2}\int_{\rm BZ}\sum_{n}f_{n}\partial_{\nu}g^{n}_{\mu\lambda}$ [77, 78]
Static susceptibility $\chi_{\mu\nu}$ $\ell^{2}/\mathcal{E}$ $\frac{e^{2}}{2\hbar}\int_{\rm BZ}\sum_{n}f_{n}G^{n}_{\mu\nu}$ [41]